### v1

# A Smooshed BMOCZ Zero Constellation for CFO Estimation Without Channel Coding

Abstract
In this study, we propose a new binary modulation on conjugate-reciprocal zeros (BMOCZ) zero constellation, which we call smooshed binary modulation on conjugate-reciprocal zeros (SBMOCZ), to address carrier frequency offset (CFO)-induced zero rotation without depending on channel coding. In our approach, we modify the phase mapping of Huffman BMOCZ by shrinking the angle between adjacent zeros, except for the first and last, to introduce a gap in the zero constellation. By discerning the gap location in the received polynomial, the receiver can estimate and correct the phase rotation. We demonstrate the error rate performance of SBMOCZ relative to Huffman BMOCZ, showing that SBMOCZ addresses a CFO-induced rotation at the cost of a modest performance reduction compared to Huffman BMOCZ in the absence of a CFO. Finally, we compare SBMOCZ to Huffman BMOCZ using a cyclically permutable code (CPC), showing a $\mathbf{4}$ dB bit error rate (BER) improvement in a fading channel, while demonstrating comparable performance across other simulations.
Index Terms:
BMOCZ, CFO, Huffman sequences, zeros of polynomials
I Introduction
Non-coherent communication schemes have emerged as a promising means to simplify receiver design and lower power consumption for ultra-massive connectivity in wireless networks [1]. Specifically, non-coherent schemes offer unique advantages in communication networks by eliminating the need for explicit CSI, which makes them appropriate for low data rate applications [2]. The lack of overhead in non-coherent schemes makes them especially beneficial in dynamic environments where CSI estimation may be impractical [3]. Since non-coherent schemes suffer from degraded performance relative to coherent ones, improving their reliability and spectral efficiency remains an active area of research.
In [4], the authors propose a new non-coherent scheme called BMOCZ, where information bits are modulated onto the zeros of the baseband signal’s $z$-transform. The zeros are constrained to lie on one of two concentric circles in the complex plane, forming a zero constellation that the authors term Huffman BMOCZ. Prior studies have explored the optimal radius for BMOCZ and examined its integration into orthogonal frequency division multiplexing (OFDM) systems [5]. The primary advantage of BMOCZ is that the information zeros are preserved regardless of the channel realization, thereby making it non-coherent. This makes BMOCZ ideal for applications requiring ultra-reliable, low-latency communication, such as intermittent short-packet transmissions in Internet of Things (IoT) networks [6]. One notable use is over-the-air computation, a technique that struggles with channel-induced distortions and requires CSI estimation, which can be impractical in dynamic environments [7]. Additionally, the desirable auto-correlation properties of BMOCZ make it an ideal choice for integrated sensing and communication [8]. In general, BMOCZ offers specific advantages as a non-coherent scheme across a wide range of potential applications.
When using BMOCZ for wireless communications, there are two notable impairments that can degrade performance: a time offset (TO) and a CFO [6]. A TO occurs whenever the start time of the transmitted signal misaligns with its actual arrival, which can lead to intersymbol interference (ISI). The authors in [6] exploit the auto-correlation properties of Huffman BMOCZ to estimate the TO and effective delay spread. The other impairment, a CFO, can significantly degrade performance. To address this problem in BMOCZ communication systems, the authors in [6] propose an affine cyclically permutable code (ACPC) combined with an oversampled direct zero-testing (DiZeT) decoder to estimate and correct the CFO. However, ACPC limits the code structure to CPCs. Therefore, to improve flexibility, alternative CFO correction methods for BMOCZ must be explored.
In this study, we propose a new zero constellation called SBMOCZ, where the angular separation between adjacent zeros, excluding the first and last, is decreased. This creates a distinct gap in the zero constellation, which will rotate under a CFO. By identifying the gap location through the received polynomial, we can correct the CFO without channel coding. We note that our algorithm for estimating a CFO-induced rotation can be implemented via a single $N$-point discrete Fourier transform (DFT), making the implementation relatively efficient. Simulation results for BER and block error rate (BLER) in additive white Gaussian noise (AWGN) and fading channels show that SBMOCZ functions under a CFO without any coding, while having a modest performance loss compared to conventional BMOCZ in the absence of CFO. Furthermore, comparisons of coded SBMOCZ and BMOCZ using ACPC show that SBMOCZ achieves a $4$ dB BER gain in a fading channel while maintaining similar performance in other scenarios.
Notation: The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$, respectively. The complex conjugate of $z=a+jb$ is expressed as $z^{\ast}=a-jb$. We denote the Euclidean norm of a vector $\boldsymbol{\rm{v}}\in\mathbb{C}^{N\times 1}$ as $||\boldsymbol{\rm{v}}||_{2}=\sqrt{\boldsymbol{\rm{v}}^{\text{H}}\boldsymbol{\rm{v}}}$. The circularly symmetric complex normal distribution with mean zero and variance $\sigma^{2}$ is expressed as $\mathcal{CN}(0,\sigma^{2})$. The uniform distribution on the interval $\left[a,b\right)$ is denoted by $\pazocal{U}_{[a,b)}$. We define $[N]=\{0,1,\dots,N-1\}$ to represent the set of the first $N$ non-negative integers. We represent the floor of a real number $a$ as $\lfloor a\rfloor$.
II System model
II-A BMOCZ Fundamentals
For a BMOCZ constellation, the $k$th bit in a message $\boldsymbol{\rm{m}}=(m_{0},m_{1},\ldots,m_{K-1})\in\{0,1\}^{K}$ is mapped to the $k$th zero of a polynomial according to
$\alpha_{k}=\begin{cases}R_{k}\,e^{j\phi_{k}},&m_{k}=1\\ R_{k}^{-1}\,e^{j\phi_{k}},&m_{k}=0\end{cases}\>,$ (1)
for $k\in[K]$, $R_{k}>1$, and $\phi_{k}\in\left[0,2\pi\right)$. By the fundamental theorem of algebra, the zeros $\boldsymbol{\alpha}=(\alpha_{0},\alpha_{1},\dots,\alpha_{K-1})\in\mathbb{C}^{K}$ uniquely define the $K$th-degree polynomial
$X(z)=\sum_{k=0}^{K}x_{k}z^{k}=x_{K}\prod_{k=0}^{K-1}(z-\alpha_{k})\>,$ (2)
where $x_{K}\neq 0$. The polynomial coefficients, given by $\boldsymbol{\rm{x}}=(x_{0},x_{1},\dots,x_{K})\in\mathbb{C}^{K+1}$, are normalized such that their squared $L_{2}$-norm satisfies $||\boldsymbol{\rm{x}}||_{2}^{2}=K+1$. Let $W(z)=\sum_{n=0}^{N_{\rm z}-1}w_{n}z^{n}$ and $H(z)=\sum_{l=0}^{L-1}h_{l}z^{l}$ denote the $z$-domain representations of the noise sequence $\boldsymbol{\rm{w}}=(w_{0},w_{1},\dots,w_{N_{\rm z}-1})\in\mathbb{C}^{N_{\rm z}}$ and the $L$-tap channel impulse response $\boldsymbol{\rm{h}}=(h_{0},h_{1},\dots,h_{L-1})\in\mathbb{C}^{L}$, respectively. Assuming transmission through an LTI channel and applying the convolution theorem for $N_{\rm z}=K+L$, the received sequence can be expressed in the $z$-domain as
$Y(z)=\sum_{n=0}^{N_{\rm z}-1}y_{n}z^{n}=X(z)H(z)+W(z)\;,$ (3)
where $\boldsymbol{\rm{y}}=(y_{0},y_{1},\dots,y_{N_{\rm z}-1})\in\mathbb{C}^{N}$ is the vector containing the coefficients of $Y(z)$. Although the presence of $H(z)$ introduces $L-1$ additional zeros to $Y(z)$, it does not alter the zeros corresponding to the message in $X(z)$. This property allows BMOCZ to operate without CSI, hence making it non-coherent. In this study, we assume a flat-fading channel where $L=1$, which means $N_{\rm z}=K+1$. This can achieved, for instance, through a time-frequency mapping of the BMOCZ polynomial coefficients in OFDM [5], which ensures $X(z)$ and $Y(z)$ have the same number of zeros.
The authors in [4] introduce Huffman BMOCZ, where all zeros are positioned on one of two concentric circles. This results in the zero mapping rule given in (1), with $R_{k}=r_{\text{hb}}$ and $\phi_{k}\triangleq 2\pi k/K$. In this scheme, $X(z)$ is called a Huffman polynomial, since the coefficients form a Huffman sequence for any combination of zeros [9, 10]. Huffman polynomials are well-conditioned, meaning that small variations in the coefficients lead to slight changes in the zeros, making them ideal for communication systems. The authors in [4] derive a simple decoding rule for BMOCZ called DiZeT, which evaluates $Y(z)$ at each conjugate-reciprocal zero pair $\mathscr{Z}_{k}=\{\alpha_{k},1/\alpha_{k}^{\ast}\}$. Using this method, the $k$th detected bit is given by
$\hat{m}_{k}=\begin{cases}1,&|Y(R_{k}\,e^{j\phi_{k}})|<R_{k}^{N_{\rm z}-1}|Y(R_{k}^{-1}\,e^{j\phi_{k}})|\\ 0,&\text{otherwise}\end{cases}\>.$ (4)
To control the minimum pairwise separation between zeros, the radius is defined as
$r_{\text{hb}}\triangleq\sqrt{1+2\lambda\sin{\left(\pi/K\right)}}\>.$ (5)
According to [4], radial zero separation has a stronger influence on BER than angular separation. Therefore, the parameter $\lambda$ is introduced as a trade-off factor to balance radial versus angular zero separation.
II-B CFO Impairment
A CFO occurs due to frequency mismatches between the transmitter and receiver oscillators. The presence of a CFO is a significant concern because it degrades the system performance. Let $\psi\in\left[0,2\pi\right)$ denote the amount of phase rotation caused by the CFO. [^1] Although this range of CFO-induced rotation is unrealistic in practice, and could disrupt the subcarrier orthogonality in OFDM, we consider it here for consistency with [6]. Due to this impairment, $Y(z)$ experiences a transformation in the $z$-domain
$\tilde{Y}(z)=\sum_{n=0}^{N_{\rm z}-1}y_{n}e^{j\psi n}z^{n}=Y(e^{j\psi}z)\>.$ (6)
Hence, all the zeros of $Y(z)$ rotate clockwise by the angle $\psi$. For a Huffman BMOCZ constellation with small $K$, this issue is less significant. However, for moderate to large values of $K$, even a small rotation causes decoding failure. Huffman BMOCZ, in the absence of channel coding, can only correct a fractional CFO due to the constant phase separation $\Delta{\phi}=2\pi/K$ between adjacent zeros and the uniform radii. Therefore, an angular rotation $\psi$ is only unique modulo $\Delta{\phi}$. Decomposing $\psi$ into a fractional component $\psi_{0}$ and an integer multiple of $\Delta{\phi}$, we obtain
$\psi=\psi_{0}+m\Delta{\phi}\;\;\text{where}\;\;\psi_{0}\in\left[0,\Delta{\phi}\right)\>,$ (7)
for any $m\in[K]$. Hence, only the fractional component $\psi_{0}$, relative to the angular separation $m\Delta{\phi}$, is detectable with uncoded Huffman BMOCZ. The authors in [6] propose using an ACPC combined with an oversampled DiZeT decoder to estimate and correct the total angular rotation. Within this scheme, the oversampled DiZeT decoder estimates the fractional component $\psi_{0}$, while the ACPC identifies the cyclic shift $m\Delta{\phi}$.
We highlight some limitations associated with ACPC. First, we note that the complexity of the code construction for ACPC increases when the code length is not a Mersenne prime, i.e., a prime number of the form $12^{K}-$ 1 for prime $K$. Furthermore, the approach requires cyclic codes, which confines the range of available coding structures. For instance, integrating BMOCZ with low-density parity check (LDPC) or polar codes is impractical within the ACPC framework. Finally, the approach introduces extra computational overhead in the decoding of the chosen cyclic code, which is required for CFO correction. For a detailed overview of ACPC, see the discussions in [6]. In this work, to retain flexibility, we propose a new BMOCZ constellation to address CFO-induced rotation without requiring any channel coding.
III Methodology
The following subsections introduce a method to estimate and correct the CFO by using a smooshed BMOCZ zero constellation.
III-A Smooshed Zero Constellation
We introduce a smooshed BMOCZ zero constellation where the phase difference between adjacent zeros, excluding the first and last, is given by $\Delta{\phi}=(2\pi-\zeta)/K$. The parameter $\zeta\in\left[0,2\pi\right)$, referred to as the smooshing factor, compresses the phase separation, which creates a larger gap between $\alpha_{0}$ and $\alpha_{K-1}$. The evaluation $|X(z)|$ along the unit circle is maximized closest to the center of the gap, since the density of zeros is lower in that region. As the constellation rotates, the gap shifts, which causes the maximum to move accordingly. Therefore, by evaluating the rotated polynomial $\tilde{Y}(z)$ at $z=e^{j\theta}$ for $\theta\in[0,2\pi)$, we can estimate $\psi$ from the $\theta$ that maximizes $|\tilde{Y}(e^{j\theta})|$. A detailed explanation for choice of the unit circle, along with a discussion of the algorithm for CFO estimation and correction, is presented in Section III-B.
We choose to center the gap on the positive real axis, which yields a new phase mapping
$\phi_{k}\triangleq\frac{(2\pi-\zeta)k}{K}+\frac{2\pi+\zeta\left(K-1\right)}{2K}\>.$ (8)
The choice of gap size affects the reliability of the CFO estimate and the displacement of zeros under noise. A small $\zeta$ results in poor CFO correction capabilities, while an excessively large $\zeta$ leads to significant zero perturbation under noise. Consequently, tuning the smooshing factor is crucial to obtain optimal performance. The choice of $\zeta$ also affects the minimum pairwise separation of the zeros, and hence the selection of radius. We modify (5) and obtain
$r_{\text{sb}}\triangleq\sqrt{1+2\lambda\sin{\left(\frac{2\pi-\zeta}{2K}\right)}}\>.$ (9)
A derivation for (9) is given in Appendix A.
When $\zeta=0$, observe that (8) and (9) reduce to Huffman BMOCZ with a $\pi/K$ counterclockwise rotation of the zeros. In this study, for simplicity, we treat SBMOCZ with $\zeta=0$ and Huffman BMOCZ interchangeably. Fig. 1 illustrates the zero constellations for Huffman BMOCZ and SBMOCZ with $K=16$, where $\zeta=0.5$ for SBMOCZ. [^2] While $\zeta$ is typically much smaller in practice, we increase it here to emphasize the zero smooshing effect. The radii are computed by setting $\lambda=0.5$ and applying (5) and (9) to derive $R_{k}$ for Huffman BMOCZ and SBMOCZ, respectively.
(a) Huffman BMOCZ.
(b) SBMOCZ with $\zeta=0.5$.
Figure 1: Comparison of Huffman BMOCZ and SBMOCZ for $K=16$. The square markers indicate possible zero locations for the transmitted polynomial $X(z)$.
(a) Huffman BMOCZ
(b) SBMOCZ with $\zeta=3{\times}10^{-2}$
(c) SBMOCZ with $\zeta=5{\times}10^{-2}$
Figure 2: $|X(z)|^{2}$ evaluated on the unit circle for $K=32$.
III-B CFO Correction with SBMOCZ
In this subsection, we introduce a method to estimate the CFO by evaluating the rotated polynomial $\tilde{Y}(z)$ at various points on the unit circle. To justify evaluating on the unit circle, we highlight some useful properties. We begin with the following proposition:
Proposition 1.
Let $X(z)$ be a polynomial with coefficients $\boldsymbol{\rm{x}}=(x_{0},x_{1},\dots,x_{K})\in\mathbb{C}^{K+1}$. Then, the squared magnitude of $X(z)$ evaluated at $z=e^{-j\theta}$ for $\theta\in\left[0,2\pi\right)$ corresponds to the discrete-time Fourier transform (DTFT) of its auto-correlation sequence $\mathbb{a}=(a_{-K},a_{-K+1},\dots,a_{K})\in\mathbb{C}^{2K+1}$. Specifically, we have
$\left|X(e^{-j\theta})\right|^{2}=\sum_{\ell=-K}^{K}a_{\ell}e^{-j\theta\ell}\;.$ (10)
This result follows from the fact that evaluating a polynomial on the unit circle yields its DTFT, and that the squared magnitude corresponds to the power spectral density, which is the DTFT of the auto-correlation sequence.
Corollary 1.
The auto-correlation sequence $\mathbb{a}$ is the same for any transmitted BMOCZ sequence $\boldsymbol{\rm{x}}$ [4]. Consequently, the evaluation $|X(e^{-j\theta})|^{2}$ yields the same result for any SBMOCZ polynomial.
The important takeaway from Proposition 1 and Corollary 1 is that the zero constellation gap will induce a peak on the unit circle at the same location for any transmitted SBMOCZ sequence $\boldsymbol{\rm{x}}$. Additionally, evaluation on the unit circle will be symmetric about the peak, as highlighted by the following lemma:
Lemma 1.
Let $X(z)$ be any SBMOCZ polynomial. When evaluating $|X(e^{-j\theta})|^{2}$ for $\theta\in[0,2\pi)$, we have
$\left|X(e^{-j\theta})\right|^{2}=\left|X(e^{j\theta})\right|^{2}\;.$ (11)
The proof is given in Appendix B. For SBMOCZ, the gap is centered on the positive real axis, resulting in a peak on the unit circle at $z=1$. To illustrate this behavior, we now present Fig. 2, which plots $|X(z)|^{2}$ along the unit circle for different $\zeta$ with $K=32$. As the gap increases in size, the peak becomes more pronounced, which makes the CFO estimate more robust against noise. However, this comes at the cost of larger zero-perturbation under noise. Additionally, Fig. 2 shows that Huffman BMOCZ is incompatible with our approach, since there is not a unique maximum on $\left[0,2\pi\right)$. Instead, evaluation on the unit circle exhibits sinusoidal behavior for Huffman BMOCZ. The key observation is that by identifying the peak location for $|\tilde{Y}(z)|^{2}$ on the unit circle, we can obtain an estimate for the angular rotation $\psi$.
Without loss of generality, we will now consider $|\tilde{Y}(z)|$, which peaks at the same location as its square. Under rotation, the maximum of $|X(e^{j(\theta+\psi)})|$ in SBMOCZ occurs when $\theta=-\psi$, so we can estimate the angular rotation from $\tilde{Y}(z)$ via
$\hat{\psi}=\arg\!\!\!\,\max_{\theta\in\left[0,2\pi\right)}\left|\tilde{Y}(e^{-j\theta})\right|\;.$ (12)
To discretize this process, we approximate $\hat{\psi}$ as
$\hat{\psi}\approx\frac{2\pi}{N}\arg\max_{n\in[N]}\left|\tilde{Y}(e^{-j2\pi n/N})\right|\>,$ (13)
where $N$ determines the resolution of the search space. We highlight that $\tilde{Y}(e^{-j2\pi n/N})$ can be evaluated using a single $N$-point DFT of the coefficients, which results in a modest time complexity of $\mathcal{O}(N\log N)$ for the proposed algorithm. To compensate for the rotation, we perform the correction $\hat{Y}(z)=\tilde{Y}(e^{-j\hat{\psi}}z)$, where $\hat{Y}(z)$ is an estimate of the received polynomial without CFO impairment. The DiZeT decoder is then applied to $\hat{Y}(z)$ to recover the binary message $\boldsymbol{\rm{m}}$.
(a) Bit error rate curves.
(b) Block error rate curves.
Figure 3: Comparison of uncoded schemes for $K=128$. The SBMOCZ scheme is configured with $r_{\text{sb}}=1.0122$ and $\zeta=0.0117$, while Huffman BMOCZ uses $r_{\text{hb}}=1.0122$.
(a) Bit error rate curves.
(b) Block error rate curves.
Figure 4: Comparison of coded schemes for $K=127$. The SBMOCZ scheme is configured with $r_{\text{sb}}=1.0123$ and $\zeta=0.0130$, while Huffman BMOCZ uses $r_{\text{hb}}=1.0123$.
IV Numerical results
In this section, we compare the performance of SBMOCZ to Huffman BMOCZ in AWGN and flat-fading channels. The noise coefficients $\boldsymbol{\rm{w}}\in\mathbb{C}^{K+1}$ are drawn from $\mathcal{CN}\left(0,\sigma_{n}^{2}\right)$, while the channel coefficient $h\in\mathbb{C}$ is drawn from $\mathcal{CN}\left(0,1\right)$ in the fading channel. To account for varying noise conditions, the noise variance $\sigma_{n}^{2}$ is computed for different $E_{\rm b}/N_{0}$. For simulations with a CFO, we draw $\psi$ from a uniform distribution $\pazocal{U}_{[0,2\pi)}$ and apply the transformation in (6) to $Y(z)$. In all simulations, we randomly sample $\boldsymbol{\rm{m}}$ from the set of all $2^{K}$ possible binary messages. Furthermore, we utilize the DiZeT decoder in all simulations and set $N=2^{10}$ for (13). The smooshing factor $\zeta$ is selected via a parameter sweep, based on minimizing BER in an AWGN channel with a CFO.
IV-A Uncoded Error Rate Performance
We begin by comparing the performance of uncoded Huffman BMOCZ to uncoded SBMOCZ with $\zeta=0.0117$. For each simulation, we set $K=128$ and $\lambda=0.5$, and determine the radii using (5) for Huffman BMOCZ and (9) for SBMOCZ. In Fig. 3 LABEL:sub@subfig:uncoded_performance_ber, we plot the BER for each scheme in AWGN and fading channels. Without a CFO, SBMOCZ performs roughly $1.3$ dB worse than Huffman BMOCZ in AWGN and $0.85$ dB worse in fading. However, when a CFO is present, Huffman BMOCZ fails (as indicated by the flat red curve), while SBMOCZ still functions but loses $1.46$ dB in AWGN and $2.92$ dB in fading relative to Huffman BMOCZ without a CFO. Observe that the BER curve with a CFO for SBMOCZ starts at a higher point compared to the version without a CFO, which occurs because the CFO estimate fails at lower $E_{\rm b}/N_{0}$, resulting in a cascade of bit errors. In Fig. 3 LABEL:sub@subfig:uncoded_performance_bler, we plot the BLER for each scheme in AWGN and fading channels. We observe that SBMOCZ performs roughly the same across different CFO conditions, but is still approximately $1.5$ dB worse in AWGN and $1$ dB in fading as compared to Huffman BMOCZ without a CFO. Again, we observe that uncoded Huffman BMOCZ fails under a CFO.
IV-B Coded Error Rate Performance
For these simulations, we utilize a (127,106)-BCH code for SBMOCZ. For Huffman BMOCZ, we employ a (127,106)-BCH absent a CFO and a (127,106)-ACPC under a CFO. To ensure fair comparison, the ACPC is implemented using a BCH code as the CPC. We note that (127,106)-BCH corrects up to three bit errors, while (127,106)-ACPC corrects a maximum of two, sacrificing one bit of error correction to detect the cyclic shift $m\Delta{\phi}$. For (127,106)-ACPC, we implement an inverse discrete Fourier transform (IDFT)-based DiZeT decoder with an oversampling factor of $Q=200$ to estimate and correct the fractional component $\psi_{0}$. Further details on the implementation of the ACPC can be found in [6]. In each scheme, we set $K=127$ and $\lambda=0.5$, using $\zeta=0.0130$ for SBMOCZ. Furthermore, we determine the radii using (5) for Huffman BMOCZ and (9) for SBMOCZ.
In Fig. 4 LABEL:sub@subfig:coded_performance_ber, we plot the BER for each scheme in AWGN and fading channels. In the AWGN channel, SBMOCZ and ACPC perform similarly at higher $E_{\rm b}/N_{0}$, while ACPC performs much worse at lower $E_{\rm b}/N_{0}$. For a higher $E_{\rm b}/N_{0}$, both SBMOCZ and ACPC perform roughly $1.6$ dB worse relative to coded Huffman BMOCZ without a CFO. However, in the fading channel, SBMOCZ achieves a significant $4$ dB gain over ACPC. In Fig. 4 LABEL:sub@subfig:coded_performance_bler, we plot the BLER for each scheme in AWGN and fading channels. We find that SBMOCZ shows a loss of around $0.65$ dB in AWGN and $0.6$ dB in fading relative to ACPC. Nevertheless, SBMOCZ demonstrates a large BER gain over ACPC in the fading channel.
V Concluding Remarks
In this study, we introduce a modified BMOCZ zero constellation termed SBMOCZ in which the angular separation between adjacent zeros, excluding the first and last, is reduced. By smooshing the zeros closer together, we create a constellation gap that rotates under a CFO, which enables the receiver to estimate and correct the rotation by identifying the gap’s position. Compared to uncoded Huffman BMOCZ, we find that uncoded SBMOCZ works under a CFO, at the cost of a modest performance reduction absent a CFO. Against Huffman BMOCZ with ACPC, we find SBMOCZ shows a $4$ dB BER gain in a fading channel, with comparable performance in other scenarios. Future work will focus on optimization of SBMOCZ constellation parameters, such as the smooshing factor and the radii for each zero. Furthermore, we will explore alternative decoding approaches for SBMOCZ.
Appendix A Radius Derivation
In SBMOCZ, the radial seperation between a conjugate-reciprocal zero pair is given by $d_{\rm cp}=r_{\text{sb}}-r_{\text{sb}}^{-1}$. The minimum separation between consecutive zeros, expressed as a function of $\zeta$, is determined using the chord length formula, which yields $d_{\rm az}=2r_{\text{sb}}^{-1}\sin\left(\frac{2\pi-\zeta}{2K}\right)$. To maximize the Euclidean distance between next-neighbor zero pairs, we equate $d_{\rm cp}$ and $d_{\rm az}$. However, since radial separation has a greater impact on BER than angular separation [4], we reintroduce the trade-off factor $\lambda\in\left(0,1\right]$ and obtain
$d_{\rm cp}$ $=\lambda d_{\rm az}$ (1a)
$r_{\text{sb}}-r_{\text{sb}}^{-1}$ $=2\lambda r_{\text{sb}}^{-1}\sin\left(\frac{2\pi-\zeta}{2K}\right)\;.$
Solving for $r_{\text{sb}}$ gives
$r_{\text{sb}}=\sqrt{1+2\lambda\sin\left(\frac{2\pi-\zeta}{2K}\right)}\;.$ (2a)
Appendix B Proof of Lemma 1
Let $\boldsymbol{\alpha}=(r_{\text{sb}}e^{j\phi_{0}},r_{\text{sb}}e^{j\phi_{1}},\dots,r_{\text{sb}}e^{j\phi_{K-1}})\in\mathbb{C}^{K}$ be the zeros of an SBMOCZ polynomial $X(z)$. Evaluating on the unit circle and taking the squared magnitude yields
$|X(e^{-j\theta})|^{2}=|x_{K}|^{2}\prod_{k=0}^{K-1}\left|e^{-j\theta}-r_{\text{sb}}e^{j\phi_{k}}\right|^{2}$ (1b)
$=|x_{K}|^{2}\prod_{k=0}^{K-1}f(\theta,\phi_{k})\;.$
Using Euler’s identities, we have
$f(\theta,\phi_{k})=\left|e^{-j\theta}-r_{\text{sb}}e^{j\phi_{k}}\right|^{2}=1+r_{\text{sb}}^{2}-2r_{\text{sb}}\cos(\theta+\phi_{k})\;.$ (2b)
Consider a case where $K$ is even. Since an SBMOCZ constellation is symmetric about the real axis, we can rewrite this product as
$|X(e^{-j\theta})|^{2}=|x_{K}|^{2}\prod_{k=0}^{\frac{K}{2}-1}f(\theta,\phi_{k})\prod_{k=0}^{\frac{K}{2}-1}f(\theta,-\phi_{k})\;.$ (3b)
Observe that negating $\theta$ does not change each product term in (3b), i.e., $f(\theta,\phi_{k})f(\theta,-\phi_{k})=f(-\theta,\phi_{k})f(-\theta,-\phi_{k})$. Therefore, we have
$\left|X(e^{-j\theta})\right|^{2}=\left|X(e^{j\theta})\right|^{2}\;.$ (4b)
When $K$ is odd, the symmetry in SBMOCZ still holds, with a single zero lying on the negative real axis. In this case, we can express (1b) as
$|X(e^{-j\theta})|^{2}=|x_{K}|^{2}f(\theta,\pi)\prod_{k=0}^{\lfloor\frac{K}{2}\rfloor-1}f(\theta,\phi_{k})\prod_{k=0}^{\lfloor\frac{K}{2}\rfloor-1}f(\theta,-\phi_{k})\;.$ (5b)
Observing that the term $\cos(\theta+\pi)=-\cos(\theta)$ in (2b) entails $f(\theta,\pi)=f(-\theta,\pi)$, we again arrive at (4b). By Corollary 1, it follows that (4b) holds for any SBMOCZ polynomial $X(z)$.