### v3

# A Smooshed BMOCZ Zero Constellation for CFO Estimation Without Channel Coding

Abstract
In this study, we propose a new binary modulation on conjugate-reciprocal zeros (BMOCZ) zero constellation, which we call smooshed binary modulation on conjugate-reciprocal zeros (SBMOCZ), to address carrier frequency offset (CFO)-induced zero rotation without depending on channel coding. In our approach, we modify the phase mapping of Huffman BMOCZ by shrinking the angle between adjacent zeros, except for the first and last, to introduce a gap in the zero constellation. By discerning the gap location in the received polynomial, the receiver can estimate and correct the phase rotation. We demonstrate the error rate performance of SBMOCZ relative to Huffman BMOCZ, showing that SBMOCZ addresses a CFO-induced rotation at the cost of a modest performance reduction compared to Huffman BMOCZ in the absence of a CFO. Finally, we compare SBMOCZ to Huffman BMOCZ using a cyclically permutable code (CPC), showing a $\mathbf{4}$ dB bit error rate (BER) improvement in a fading channel, while demonstrating comparable performance across other simulations.
Index Terms:
BMOCZ, CFO, Huffman sequences, zeros of polynomials
I Introduction
Non-coherent communication has emerged as a promising means to simplify receiver design and lower power consumption for ultra-massive connectivity in wireless networks[1]. In particular, non-coherent communication offers unique advantages in wireless networks by eliminating the need for explicit CSI, which makes it suitable for low data rate applications[2]. The lack of overhead in non-coherent communication makes it especially beneficial in dynamic environments where CSI estimation may be impractical[3]. Since non-coherent communication suffers from degraded performance compared to coherent communication, improving its reliability and spectral efficiency remains an active area of research.
In [4], the authors propose a novel non-coherent modulation scheme called BMOCZ, where the information bits are mapped to the zeros of the baseband signal’s $z$-transform. The information zeros are constrained to lie on one of two concentric circles in the complex plane, forming a zero constellation which the authors term Huffman BMOCZ. Previous studies have explored the optimal radius for BMOCZ, and examined its integration into orthogonal frequency division multiplexing (OFDM)[5]. The primary advantage of BMOCZ is that the information zeros are preserved at the receiver regardless of the channel realization, thereby making it non-coherent. This makes BMOCZ ideal for applications requiring ultra-reliable, low-latency communication, such as intermittent short-packet transmissions in Internet of Things (IoT) networks[6]. One notable use is over-the-air computation, a technique that struggles with channel-induced distortions and requires CSI estimation, which can be impractical in dynamic environments [7]. Moreover, the desirable auto-correlation properties of BMOCZ make it an ideal choice for integrated sensing and communication[8]. In general, BMOCZ offers specific advantages as a non-coherent modulation scheme across a wide range of potential applications.
When utilizing BMOCZ for wireless communication, there are two notable impairments that can degrade the error rate performance: a time offset (TO) and CFO [6]. A TO occurs whenever the start time of the transmitted signal misaligns with its actual arrival, leading to intersymbol interference (ISI). In [6], the authors exploit the auto-correlation properties of Huffman BMOCZ to estimate and correct the TO. The CFO can also significantly degrade the error rate performance. To address the CFO in BMOCZ-based communication, the authors in [6] introduce an affine cyclically permutable code (ACPC) combined with an oversampled direct zero-testing (DiZeT) decoder to estimate and correct the CFO. However, the ACPC limits the code structure to CPCs. Therefore, to improve flexibility, alternative CFO correction methods for BMOCZ must be explored.
In this study, we propose a new smooshed zero constellation called SBMOCZ, where the angular separation between adjacent zeros, excluding the first and last, is decreased. This creates a distinct gap in the zero constellation, which will rotate under a CFO. By identifying the gap location in the received polynomial, we can correct the CFO without channel coding. We highlight that our algorithm for estimating a CFO-induced rotation can be implemented via a single $N$-point discrete Fourier transform (DFT), making the implementation relatively efficient. The simulation results for BER and block error rate (BLER) in additive white Gaussian noise (AWGN) and fading channels demonstrate that SBMOCZ functions under a CFO without any channel coding, while incurring a modest performance loss compared to Huffman BMOCZ without a CFO. Furthermore, comparisons of coded SBMOCZ and Huffman BMOCZ with an ACPC show that SBMOCZ achieves a $4$ dB BER gain in the fading channel, while maintaining similar performance across other scenarios.
Notation: The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$, respectively. We define $\mathbb{Z}^{N}_{2}$ as the set of all binary vectors of length $N$. The imaginary unit is given by ${\rm j}=\sqrt{-1}$, and Euler’s number is represented with ${\rm e}$. The complex conjugate of $z=a+{\rm j}b$ is expressed as $z^{\ast}=a-{\rm j}b$. We denote the Euclidean norm of a vector $\boldsymbol{\rm{v}}\in\mathbb{C}^{N\times 1}$ as $||\boldsymbol{\rm{v}}||_{2}=\sqrt{\boldsymbol{\rm{v}}^{\text{H}}\boldsymbol{\rm{v}}}$. The circularly symmetric complex normal distribution with mean zero and variance $\sigma^{2}$ is expressed as $\mathcal{CN}(0,\sigma^{2})$. The uniform distribution on the interval $\left[a,b\right)$ is denoted by $\pazocal{U}_{[a,b)}$. We define $[N]=\{0,1,\dots,N-1\}$ to represent the set of the first $N$ non-negative integers. Finally, we express the floor of a real number $a$ as $\lfloor a\rfloor$.
II System model
II-A BMOCZ Fundamentals
[…]
$\alpha_{k}=\begin{cases}R_{k}\,{\rm e}^{{\rm j}\phi_{k}},&m_{k}=1\\ R_{k}^{-1}\,{\rm e}^{{\rm j}\phi_{k}},&m_{k}=0\end{cases}\>,\ \ k\in[K]\;,$ (1)
with $R_{k}>1$, and $\phi_{k}\in\left[0,2\pi\right)$. By the fundamental theorem of algebra, the zeros $\boldsymbol{\alpha}=[\alpha_{0},\alpha_{1},\dots,\alpha_{K-1}]\in\mathbb{C}^{K}$ uniquely define the $K$th degree polynomial
$X(z)=\sum_{k=0}^{K}x_{k}z^{k}=x_{K}\prod_{k=0}^{K-1}(z-\alpha_{k})\;,$ (2)
where $x_{K}\neq 0$. The polynomial coefficients, given by $\boldsymbol{\rm{x}}=[x_{0},x_{1},\dots,x_{K}]\in\mathbb{C}^{K+1}$, are normalized such that $||\boldsymbol{\rm{x}}||_{2}^{2}=K+1$. Let $W(z)=\sum_{n=0}^{N_{\rm z}-1}w_{n}z^{n}$ and $H(z)=\sum_{l=0}^{L-1}h_{l}z^{l}$ indicate the $z$-domain representations for the noise sequence $\boldsymbol{\rm{w}}=[w_{0},w_{1},\dots,w_{N_{\rm z}-1}]\in\mathbb{C}^{N_{\rm z}}$ and the $L$-tap channel impulse response $\boldsymbol{\rm{h}}=[h_{0},h_{1},\dots,h_{L-1}]\in\mathbb{C}^{L}$, respectively. Assuming transmission through an LTI channel and applying the convolution theorem for $N_{\rm z}=K+L$, the received sequence can be expressed in the $z$-domain as
$Y(z)=\sum_{n=0}^{N_{\rm z}-1}y_{n}z^{n}=X(z)H(z)+W(z)\;,$ (3)
where $\boldsymbol{\rm{y}}=[y_{0},y_{1},\dots,y_{N_{\rm z}-1}]\in\mathbb{C}^{N_{\rm z}}$ is the vector containing the coefficients of $Y(z)$. Although the presence of $H(z)$ introduces $L-1$ additional zeros to $Y(z)$, it does not alter the zeros corresponding to the message in $X(z)$. This property allows BMOCZ to operate without CSI, thereby making it non-coherent. In this study, we assume a flat-fading channel where $L=1$, which means $N_{\rm z}=K+1$. This can be achieved, for example, through a time-frequency mapping of BMOCZ polynomial coefficients in OFDM, which ensures $X(z)$ and $Y(z)$ have the same number of zeros [5].
In [4], the authors introduce Huffman BMOCZ, where the information zeros are positioned on one of two concentric circles in the complex plane. This results in the zero mapping rule provided in (1), with $R_{k}=r_{\text{hb}}$ and $\phi_{k}\triangleq 2\pi k/K$. In this scheme, $X(z)$ is called a Huffman polynomial, since the coefficients form a Huffman sequence for any combination of information zeros [9, 10]. Huffman polynomials are well-conditioned, meaning that small variations in the coefficients lead to slight changes in the zeros, making them ideal for communication systems. In [4], the authors derive a simple decoding rule for BMOCZ called DiZeT, which evaluates $Y(z)$ at each conjugate-reciprocal zero pair $\mathscr{Z}_{k}\triangleq\{\alpha_{k},1/\alpha_{k}^{\ast}\}$. Using this method, the $k$th detected bit is given by
$\hat{m}_{k}=\begin{cases}1,&\left|Y(R_{k}\,{\rm e}^{{\rm j}\phi_{k}})\right|<R_{k}^{N_{\rm z}-1}\left|Y(R_{k}^{-1}\,{\rm e}^{{\rm j}\phi_{k}})\right|\\ 0,&\text{otherwise}\end{cases}\;.$ (4)
To control the minimum pairwise separation between zeros, the radius is defined as
$r_{\text{hb}}\triangleq\sqrt{1+2\lambda\sin{\left(\pi/K\right)}}\>.$ (5)
According to [4], radial zero separation has a stronger influence on BER than angular separation. Therefore, the parameter $\lambda\in\left(0,1\right]$ is introduced as a trade-off factor to balance radial versus angular zero separation.
II-B CFO Impairment
A CFO occurs due to frequency mismatches between the transmitter and receiver oscillators. The presence of a CFO in Huffman BMOCZ is a significant concern because it degrades the system performance. Let $\psi\in\left[0,2\pi\right)$ denote the amount of phase rotation caused by the CFO. [^1] Although this range of CFO-induced rotation is unrealistic in practice, and could disrupt the subcarrier orthogonality in OFDM, we consider it here for consistency with [6]. Due to this impairment, $Y(z)$ experiences a transformation in the $z$-domain
$\tilde{Y}(z)=\sum_{n=0}^{N_{\rm z}-1}y_{n}{\rm e}^{{\rm j}\psi n}z^{n}=Y(z{\rm e}^{{\rm j}\psi})\;.$ (6)
Hence, the zeros of $Y(z)$ rotate clockwise by the angle $\psi$. For a Huffman BMOCZ constellation with small $K$, this issue is less significant. However, for moderate or large values of $K$, even a small rotation causes decoding failure. Huffman BMOCZ, in the absence of channel coding, can only correct a fractional CFO due to the constant phase separation $\Delta{\phi}=2\pi/K$ between adjacent zeros and the uniform radii. Therefore, an angular rotation $\psi$ is only unique modulo $\Delta{\phi}$. Decomposing $\psi$ into a fractional component $\psi_{0}$ and an integer multiple of $\Delta{\phi}$, we obtain
$\psi=\psi_{0}+m\Delta{\phi}\;\;\text{where}\;\;\psi_{0}\in\left[0,\Delta{\phi}\right)\;,$ (7)
for any $m\in[K]$. Therefore, only the fractional component $\psi_{0}$, relative to the angular separation $m\Delta{\phi}$, is detectable with uncoded Huffman BMOCZ. To address this issue, the authors in [6] propose an ACPC combined with an oversampled DiZeT decoder to estimate and correct the total zero rotation. Within this framework, the oversampled DiZeT decoder estimates the fractional component $\psi_{0}$, while the ACPC identifies the cyclic shift $m\Delta{\phi}$.
We highlight several limitations associated with the ACPC. First, we note that the complexity of the code construction for an ACPC increases when the code length is not a Mersenne prime, i.e., a prime number of the form $12^{K}-$ 1 for prime $K$. Furthermore, the approach requires using cyclic codes, which confines the range of available coding structures. For instance, integrating BMOCZ with low-density parity check (LDPC) or polar codes is not practical within the ACPC framework. Finally, the approach introduces extra computational overhead in the decoding of the chosen cyclic code, which is required to correct the zero rotation. For a detailed overview of ACPC, we direct the reader to [6]. In this work, to retain flexibility, we propose a new BMOCZ zero constellation to address CFO-induced rotation without any channel coding.
III Methodology
The following subsections introduce a framework to estimate and correct CFO-induced rotation by using a smooshed BMOCZ zero constellation.
III-A Smooshed Zero Constellation
We introduce a smooshed BMOCZ zero constellation where the phase difference between adjacent zeros, excluding the first and last, is given by $\Delta{\phi}=(2\pi-\zeta)/K$. The parameter $\zeta\in\left[0,2\pi\right)$, referred to as the smooshing factor, compresses the phase separation, which creates a larger gap between $\alpha_{0}$ and $\alpha_{K-1}$. The evaluation of $|X(z)|$ along the unit circle is maximized closest to the center of the gap, since the density of zeros is lower in that region. As the constellation rotates, the gap shifts, which causes the maximum to move accordingly. Therefore, by evaluating the rotated polynomial $\tilde{Y}(z)$ at $z={\rm e}^{-{\rm j}\theta}$ for $\theta\in[0,2\pi)$, we can estimate $\psi$ from the $\theta$ that maximizes $|\tilde{Y}({\rm e}^{-{\rm j}\theta})|$. A detailed explanation for choosing the unit circle, along with a discussion of the algorithm for CFO estimation and correction, is provided in Section III-B.
We choose to center the gap on the positive real axis, which yields a new phase mapping
$\phi_{k}\triangleq\frac{(2\pi-\zeta)k}{K}+\frac{2\pi+\zeta\left(K-1\right)}{2K}\;.$ (8)
The selection of $\zeta$ affects the reliability of CFO correction and the displacement of zeros under noise. In particular, a small $\zeta$ results in poor CFO correction capabilities, while an excessively large $\zeta$ leads to significant zero perturbation under noise. Therefore, careful selection of the smooshing factor is required to obtain good error rate performance. The choice of $\zeta$ also affects the minimum pairwise separation of the zeros, and hence the selection of radius. We modify (5) and obtain
$r_{\text{sb}}\triangleq\sqrt{1+2\lambda\sin{\left(\frac{2\pi-\zeta}{2K}\right)}}\;.$ (9)
A derivation for (9) is given in Appendix A.
When $\zeta=0$, observe that (8) and (9) reduce to Huffman BMOCZ with a $\pi/K$ counterclockwise rotation of the zeros. In this study, for simplicity, we treat SBMOCZ with $\zeta=0$ and Huffman BMOCZ interchangeably. Fig. 1 illustrates the zero constellations for Huffman BMOCZ and SBMOCZ with $K=16$, where $\zeta=1/2$ for SBMOCZ. [^2] While $\zeta$ is typically much smaller in practice, we increase it here to emphasize the zero smooshing effect. The radii are computed by setting $\lambda=1/2$ and applying (5) and (9) to derive $R_{k}$ for Huffman BMOCZ and SBMOCZ, respectively.
(a) Huffman BMOCZ.
(b) SBMOCZ with $\zeta=1/2$.
Figure 1: Comparison of Huffman BMOCZ and SBMOCZ for $K=16$. The square markers indicate possible zero locations for the transmitted polynomial $X(z)$.
(a) Huffman BMOCZ.
(b) SBMOCZ with $\zeta=3/100$.
(c) SBMOCZ with $\zeta=1/20$.
Figure 2: $|X(z)|^{2}$ evaluated on the unit circle for $K=32$.
III-B CFO Correction with SBMOCZ
In this subsection, we introduce a method to correct a CFO impairment by evaluating the rotated polynomial $\tilde{Y}(z)$ at various points along the unit circle. To justify evaluating on the unit circle, we highlight some useful properties. We begin with the following proposition:
Proposition 1.
Let $X(z)$ be a polynomial with the coefficient vector $\boldsymbol{\rm{x}}=[x_{0},x_{1},\dots,x_{K}]\in\mathbb{C}^{K+1}$. Then, the squared magnitude of $X({\rm e}^{-{\rm j}\theta})$ for $\theta\in\left[0,2\pi\right)$ corresponds to the discrete-time Fourier transform (DTFT) of the auto-correlation sequence $\mathbb{a}=[a_{-K},a_{-K+1},\dots,a_{K}]\in\mathbb{C}^{2K+1}$ for $\boldsymbol{\rm{x}}$. In particular, we have
$\left|X({\rm e}^{-{\rm j}\theta})\right|^{2}=\sum_{\ell=-K}^{K}a_{\ell}{\rm e}^{-{\rm j}\theta\ell}\;.$ (10)
This result follows from the fact that evaluating a polynomial on the unit circle yields its DTFT, and that the squared magnitude corresponds to the power spectral density, which is the DTFT of the auto-correlation sequence.
Corollary 1.
The auto-correlation sequence $\mathbb{a}\in\mathbb{C}^{2K+1}$ is the same for any transmitted BMOCZ coefficient sequence $\boldsymbol{\rm{x}}\in\mathbb{C}^{K+1}$ [4]. Consequently, for any SBMOCZ polynomial $X(z)$, the evaluation $|X({\rm e}^{-{\rm j}\theta})|^{2}$ yields the same result.
The important takeaway from Proposition 1 and Corollary 1 is that the smooshed zero constellation gap will induce a peak on the unit circle at the same location for any transmitted SBMOCZ coefficient sequence $\boldsymbol{\rm{x}}$. Additionally, evaluation on the unit circle will be symmetric about the peak, as highlighted by the following lemma:
Lemma 1.
Let $X(z)$ be an SBMOCZ polynomial. When evaluating $|X({\rm e}^{-{\rm j}\theta})|^{2}$ with $\theta\in[0,2\pi)$, we have
$|X({\rm e}^{-{\rm j}\theta})|^{2}=|X({\rm e}^{{\rm j}\theta})|^{2}\;.$ (11)
The proof is given in Appendix B. For SBMOCZ, the gap is centered on the positive real axis, resulting in a peak on the unit circle at $z=1$. To illustrate this behavior, we now present Fig. 2, which plots $|X(z)|^{2}$ along the unit circle for different $\zeta$ with $K=32$. As the gap increases in size, the peak becomes more pronounced, which makes the CFO estimate more robust against noise. However, this comes at the price of larger zero perturbation under noise. Additionally, Fig. 2 shows that Huffman BMOCZ is incompatible with our approach, since there is not a unique maximum on $\left[0,2\pi\right)$. Instead, evaluation on the unit circle exhibits sinusoidal behavior for Huffman BMOCZ. The key observation is that by identifying the peak location for $|\tilde{Y}(z)|^{2}$ on the unit circle, we can obtain an estimate for the angular rotation $\psi$.
Without loss of generality, we will now consider $|\tilde{Y}(z)|$, which peaks at the same location as its square. Under rotation, the maximum of $|X({\rm e}^{{\rm j}(\theta+\psi)})|$ in SBMOCZ occurs whenever $\theta=-\psi$, so we can estimate the angular rotation from $\tilde{Y}(z)$ via
$\hat{\psi}=\arg\!\!\!\,\max_{\theta\in\left[0,2\pi\right)}\left|\tilde{Y}({\rm e}^{-{\rm j}\theta})\right|\;.$ (12)
To discretize this process, we approximate $\hat{\psi}$ as
$\hat{\psi}\approx\frac{2\pi}{N}\arg\max_{n\in[N]}\left|\tilde{Y}({\rm e}^{-{\rm j}2\pi n/N})\right|\;,$ (13)
where $N$ determines the resolution of the search space. We highlight that $\tilde{Y}({\rm e}^{-{\rm j}2\pi n/N})$ can be evaluated using a single $N$-point DFT of the coefficients, which results in a modest time complexity of $\mathcal{O}(N\log N)$ for the proposed algorithm. To compensate for the rotation, we perform the correction $\hat{Y}(z)=\tilde{Y}(z{\rm e}^{-{\rm j}\hat{\psi}})$, where $\hat{Y}(z)$ is an estimate of the received polynomial without a CFO impairment. The DiZeT decoder is then applied to $\hat{Y}(z)$ to recover the binary message $\boldsymbol{\rm{m}}\in\mathbb{Z}^{K}_{2}$.
(a) Bit error rate curves.
(b) Block error rate curves.
Figure 3: Comparison of uncoded schemes for $K=128$. The SBMOCZ scheme is configured with $r_{\text{sb}}=1.0122$ and $\zeta=0.0117$, while Huffman BMOCZ uses $r_{\text{hb}}=1.0122$.
(a) Bit error rate curves.
(b) Block error rate curves.
Figure 4: Comparison of coded schemes for $K=127$. The SBMOCZ scheme is configured with $r_{\text{sb}}=1.0123$ and $\zeta=0.0130$, while Huffman BMOCZ uses $r_{\text{hb}}=1.0123$.
IV Numerical results
In this section, we compare the performance of SBMOCZ to Huffman BMOCZ in AWGN and flat-fading channels. The noise coefficients $\boldsymbol{\rm{w}}\in\mathbb{C}^{K+1}$ are drawn from $\mathcal{CN}\left(0,\sigma_{n}^{2}\right)$, while the channel coefficient $h\in\mathbb{C}$ is drawn from $\mathcal{CN}\left(0,1\right)$ for the fading channel. To account for varying noise conditions, the noise variance $\sigma_{n}^{2}$ is computed for different $E_{\rm b}/N_{0}$. For simulations using a CFO, we draw $\psi$ from the uniform distribution $\pazocal{U}_{[0,2\pi)}$ and apply the transformation in (6) to $Y(z)$. In each simulation, we randomly sample $\boldsymbol{\rm{m}}\in\mathbb{Z}^{K}_{2}$ from the set of all $2^{K}$ possible binary messages. Furthermore, we utilize the DiZeT decoder and set $N=1,024$ for the estimator in (13). The smooshing factor $\zeta$ is selected using a parameter sweep, based on minimizing BER under a CFO in AWGN.
IV-A Uncoded Error Rate Performance
To begin, we compare the performance of uncoded Huffman BMOCZ to uncoded SBMOCZ using $\zeta=0.0117$. For each simulation, we set $K=128$ and $\lambda=1/2$, and determine the radii using (5) for Huffman BMOCZ and (9) for SBMOCZ. In Fig. 3LABEL:sub@subfig:uncoded_performance_ber, we plot the BER curves for each scheme in AWGN and fading channels. Without a CFO, SBMOCZ performs roughly $1.3$ dB worse than Huffman BMOCZ in AWGN and $0.85$ dB worse in fading. However, when a CFO is introduced, Huffman BMOCZ fails, as indicated by the red curve, while SBMOCZ remains functional but loses $1.46$ dB in AWGN and $2.92$ dB in fading relative to Huffman BMOCZ without a CFO. Observe that the BER curve for SBMOCZ with a CFO starts at a higher point compared to the version without a CFO, which occurs because the CFO estimate fails at lower $E_{\rm b}/N_{0}$, resulting in a cascade of bit errors. In Fig. 3LABEL:sub@subfig:uncoded_performance_bler, we plot the BLER for each scheme in AWGN and fading channels. We observe that SBMOCZ performs roughly the same with and without a CFO, but is still approximately $1.5$ dB worse in AWGN and $1$ dB in fading as compared to Huffman BMOCZ without a CFO. Again, we see that Huffman BMOCZ without channel coding does not function under a CFO.
IV-B Coded Error Rate Performance
For these simulations, we utilize a (127,106)-BCH code for SBMOCZ. For Huffman BMOCZ, we employ a (127,106)-BCH code absent a CFO and a (127,106)-ACPC under a CFO. To ensure fair comparison, the ACPC is implemented using a BCH code as the CPC. We note that (127,106)-BCH corrects up to three bit errors, while (127,106)-ACPC only corrects a maximum of two, sacrificing one bit of error correction to detect the cyclic shift $m\Delta{\phi}$. For (127,106)-ACPC, we implement an inverse discrete Fourier transform (IDFT)-based DiZeT decoder with an oversampling factor of $Q=200$ to estimate and correct the fractional component $\psi_{0}$. Further details on the implementation of the ACPC can be found in [6]. In each scheme, we set $K=127$ and $\lambda=1/2$, using $\zeta=0.0130$ for SBMOCZ. Furthermore, we determine the radii using (5) for Huffman BMOCZ and (9) for SBMOCZ.
In Fig. 4LABEL:sub@subfig:coded_performance_ber, we plot the BER for each scheme in AWGN and fading channels. In AWGN, SBMOCZ and the ACPC perform similarly at higher $E_{\rm b}/N_{0}$, while the ACPC performs much worse at lower $E_{\rm b}/N_{0}$. For higher $E_{\rm b}/N_{0}$, both SBMOCZ and the ACPC perform roughly $1.6$ dB worse relative to coded Huffman BMOCZ without a CFO. However, in the fading channel, SBMOCZ achieves a significant $4$ dB gain over the ACPC. In Fig. 4LABEL:sub@subfig:coded_performance_bler, we plot the BLER for each scheme in AWGN and fading channels. We find that SBMOCZ shows a loss of around $0.65$ dB in AWGN and $0.6$ dB in fading compared to the ACPC. Nevertheless, SBMOCZ demonstrates a large BER gain over the ACPC in a fading channel.
V Concluding Remarks
In this study, we introduce a new smooshed BMOCZ zero constellation called SBMOCZ in which the angular separation between adjacent zeros, excluding the first and last, is reduced. By smooshing the zeros closer together, we create a constellation gap that rotates under a CFO, which enables the receiver to estimate and correct the rotation by identifying the gap’s position. Compared to uncoded Huffman BMOCZ, we find that uncoded SBMOCZ functions under a CFO, at the cost of a modest performance reduction without a CFO. Against Huffman BMOCZ with an ACPC, we find that coded SBMOCZ achieves a $4$ dB BER gain in the fading channel, with comparable performance in other scenarios. Future work will focus on optimization of SBMOCZ constellation parameters, such as the smooshing factor and the radii for each zero. Moreover, we will explore alternative decoding approaches for SBMOCZ.
Appendix A Radius Derivation for SBMOCZ
In SBMOCZ, the radial seperation between a conjugate-reciprocal zero pair is given by $d_{\rm cp}=r_{\text{sb}}-r_{\text{sb}}^{-1}$. The minimum separation between consecutive zeros, expressed as a function of $\zeta$, is determined using the chord length formula, which yields $d_{\rm az}=2r_{\text{sb}}^{-1}\sin\left(\frac{2\pi-\zeta}{2K}\right)$. To maximize the Euclidean distance between next-neighbor zero pairs, we equate $d_{\rm cp}$ and $d_{\rm az}$. However, since radial separation has a greater impact on BER than angular separation [4], we reintroduce the trade-off factor $\lambda\in\left(0,1\right]$ and obtain
$d_{\rm cp}$ $=\lambda d_{\rm az}$ (1a)
$r_{\text{sb}}-r_{\text{sb}}^{-1}$ $=2\lambda r_{\text{sb}}^{-1}\sin\left(\frac{2\pi-\zeta}{2K}\right)\;.$
Solving for $r_{\text{sb}}$ gives
$r_{\text{sb}}=\sqrt{1+2\lambda\sin\left(\frac{2\pi-\zeta}{2K}\right)}\;.$ (2a)
Appendix B Proof of Lemma 1
Let $\boldsymbol{\alpha}=[r_{\text{sb}}{\rm e}^{{\rm j}\phi_{0}},r_{\text{sb}}{\rm e}^{{\rm j}\phi_{1}},\dots,r_{\text{sb}}{\rm e}^{{\rm j}\phi_{K-1}}]\in\mathbb{C}^{K}$ be the zeros of an SBMOCZ polynomial $X(z)$. Evaluating on the unit circle and taking the squared magnitude yields
$|X({\rm e}^{-{\rm j}\theta})|^{2}=|x_{K}|^{2}\prod_{k=0}^{K-1}\left|{\rm e}^{-{\rm j}\theta}-r_{\text{sb}}{\rm e}^{{\rm j}\phi_{k}}\right|^{2}$ (1b)
$=|x_{K}|^{2}\prod_{k=0}^{K-1}f(\theta,\phi_{k})\;.$
Using Euler’s identities, we have
$f(\theta,\phi_{k})=\left|{\rm e}^{-{\rm j}\theta}-r_{\text{sb}}{\rm e}^{{\rm j}\phi_{k}}\right|^{2}=1+r_{\text{sb}}^{2}-2r_{\text{sb}}\cos(\theta+\phi_{k})\;.$ (2b)
Consider a case where $K$ is even. Since an SBMOCZ constellation is symmetric about the real axis, we can rewrite this product as
$|X({\rm e}^{-{\rm j}\theta})|^{2}=|x_{K}|^{2}\prod_{k=0}^{\frac{K}{2}-1}f(\theta,\phi_{k})\prod_{k=0}^{\frac{K}{2}-1}f(\theta,-\phi_{k})\;.$ (3b)
Observe that negating $\theta$ does not change each product term in (3b), i.e., $f(\theta,\phi_{k})f(\theta,-\phi_{k})=f(-\theta,\phi_{k})f(-\theta,-\phi_{k})$. Therefore, we have
$\left|X({\rm e}^{-{\rm j}\theta})\right|^{2}=\left|X({\rm e}^{{\rm j}\theta})\right|^{2}\;.$ (4b)
When $K$ is odd, the symmetry in SBMOCZ still holds, with a single zero lying on the negative real axis. In this case, we can express (1b) as
$|X({\rm e}^{-{\rm j}\theta})|^{2}=|x_{K}|^{2}f(\theta,\pi)\prod_{k=0}^{\lfloor\frac{K}{2}\rfloor-1}f(\theta,\phi_{k})\prod_{k=0}^{\lfloor\frac{K}{2}\rfloor-1}f(\theta,-\phi_{k})\;.$ (5b)
Observing that the term $\cos(\theta+\pi)=-\cos(\theta)$ in (2b) entails $f(\theta,\pi)=f(-\theta,\pi)$, we again arrive at (4b). By Corollary 1, it follows that (4b) holds for any SBMOCZ polynomial $X(z)$.